OpenStudy (anonymous):

f'(x)=x+3e^x find f and f(0)=5

OpenStudy (anonymous):

Whats ur derivitive? f'(x)= x+3e^x exponention function?

OpenStudy (anonymous):

\(\large f(x)=\int{x+3e^x}dx=\frac{1}{2}x^2+3e^x+C \)

OpenStudy (anonymous):

f(0)=5 ---> 5=3+C ---> C=2

OpenStudy (anonymous):

d/dy(e^x)=e^x so ur function will be:f(x)=1/2x^2+3e^x